OpenStudy (anonymous):

LAst Question!!! Need help!! Tracy is finding the equation of a median-median line. She found that the equation of the line through the first and third summary points is y = 0.4x + 38. Then she found that the equation of the line that is parallel to this line and passes through the second summary point is y = 0.4x + 59. What is the equation of the median-median line? a. y = 0.4x + 41 b. y = 0.4x + 45 c. y = 0.4x + 52 d. y = 0.4x + 56 Please explain! i don't get this at all :(

OpenStudy (anonymous):

Does anybody get this problem???

OpenStudy (anonymous):

first find median of both the lines than u will get two points.than use those two points to get a equation of line i think that would be the solution

OpenStudy (anonymous):

is the solution y=0.4x+45?? or is it something else? idk haha

OpenStudy (anonymous):

line between these two points would be median median line i guess

OpenStudy (apoorvk):

@Callisto - your domain!

jimthompson5910 (jim_thompson5910):

According to this page http://www.amstat.org/publications/jse/v14n2/morrell.html#Figure1 the median median line is a regression line that is found by first finding the line that passes through two medians and then you find the equation of the parallel line that passes through the third remaining median. From there, you calculate the distance between the parallel lines. Then you divide that distance into 3 (because you have 3 medians). Finally, you mark the point along the perpendicular segment that connects the two parallel lines that is exactly 1/3 of the distance away from the first line and 2/3 of the distance from the second line. Basically, we're doing a weighted average which tells us that the line that is passing through 2 medians has more "pull" so to speak.

OpenStudy (apoorvk):

@eliassaab - Sir could you please help us? :)

OpenStudy (anonymous):

NVM it's okay! thanks for all your help! :) i gotta go now i think lol :)

jimthompson5910 (jim_thompson5910):

It's definitely a lot to take in, but the good news is this: you're given the two lines at the top of the problem. This means that you simply need to find a parallel line to both equations and this new line must be between the two parallel lines. And, this new line must be positioned in such a way that the ratio of the distances from it to the first and second line is 1:2 Essentially, we're cutting the distance between the two given parallel lines in 3 parts So in the end, we're simply cutting the distance of 59-38 = 21 into 3 parts to get 21/3 = 7 Then we're adding this to the y-intercept of the first equation: 7+38 = 45 So the answer is y = 0.4x + 45